2. An aluminum wire, of length li 60.0cm, cross sectional area 1.00 x 10-2cm2, density of 2.60g/cm3, is connected to a steel wire of density 7.80g/cm3 and the same cross sectional area. The compound wire, loaded with a block of mass m 10.0kg, is arranged as shown below so that the distance l2 from the joint to the supporting pulley is 86.6cm. Transverse waves are set up in the wire by an external source of variable frequency. (a) Find the lowest frequency of excitation for which standing waves are observed such that the joint in the wire is a node. (b) How many nodes are observed at this frequency? 2An aluminum wire, of length l1 = 60.0cm, cross sectional area 1.00 × 10 cm , density waves(from(the(sun(just(outs3ide(the(atmospheres(of(Venus,(the(earth,(and(Mars.((valu3es(can(be( of 2.60g/cm , is connected to a steel wire of density 7.80g/cm and the same cross sectional found(in(the(appendix(of(your(textbook)( ( Problem(2:( area. The compound wire, loaded with a block of mass m = 10.0kg, is arranged as shown below so that the distance l2 from the joint to the supporting pulley is 86.6cm. Transverse waves are set up in the wire by an external source of variable frequency. (a) Find the lowest frequency of excitation for which standing waves are observed such that the joint in the wire is a node. (b) How many nodes are observed at this frequency?
Answer #1

Number of nodes in the wire, excluding end points are (n-1).

Nodes in Aluminium wire = 2-1 =1

Nodes in steel wire = 5-1 =4

One node is at junction. Total number of nodes are 6, excluding end point nodes.